$A$ bag contains $3$ red balls and $5$ black balls. $A$ ball is drawn at random from the bag. What is the probability that the ball drawn is
$(i)$ red?
$(ii)$ not red?

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(A) $(i)$ Total number of balls in the bag $= 3 + 5 = 8$.
The probability of drawing a red ball is given by the formula: $P(\text{red}) = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}}$.
$P(\text{red}) = \frac{3}{8}$.
$(ii)$ The probability of not drawing a red ball is the complement of drawing a red ball.
$P(\text{not red}) = 1 - P(\text{red})$.
$P(\text{not red}) = 1 - \frac{3}{8} = \frac{8 - 3}{8} = \frac{5}{8}$.

Explore More

Similar Questions

$A$ box contains $90$ discs which are numbered from $1$ to $90$. If one disc is drawn at random from the box, find the probability that it bears:
$(i)$ a two-digit number
$(ii)$ a perfect square number
$(iii)$ a number divisible by $5$.

Difficult
View Solution

Complete the following statements:
$(i)$ Probability of an event $E +$ Probability of the event 'not $E$' $= \dots \dots \dots \dots \dots$
$(ii)$ The probability of an event that cannot happen is $\dots \dots \dots$. Such an event is called $\dots \dots \dots \dots \dots$
$(iii)$ The probability of an event that is certain to happen is $\dots \dots \dots$. Such an event is called $\dots \dots \dots \dots \dots$

There are $40$ students in Class $X$ of a school,of whom $25$ are girls and $15$ are boys. The class teacher has to select one student as a class representative. She writes the name of each student on a separate card,the cards being identical. Then she puts cards in a bag and stirs them thoroughly. She then draws one card from the bag. What is the probability that the name written on the card is the name of $(i)$ a girl? $(ii)$ a boy?

$A$ missing helicopter is reported to have crashed somewhere in the rectangular region shown in the figure. What is the probability that it crashed inside the lake shown in the figure?

Harpreet tosses two different coins simultaneously (say,one is of $Rs. 1$ and other of $Rs. 2$). What is the probability that she gets at least one head?

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo